Daddyleo:

lol have a question ok so divide 6 by 3 wat u get den multiply it by 20 den sub 5 den multiply aging the number you see previously tell me if ya confused

jhonyy9:

step by step 6/3 =2 ×20 = 40 -5 = 35 what s the next ?

Xio:

I a assuming this is a puzzle. Am I wrong? I believe jhonny that what he fans is we got to multiply 35 by 40 afterwards but I'm not completely sure. Its difficult to understand because for the way he worded it.

jhonyy9:

@daddy14 your opinion pls about this ?

Daddyleo:

you forgot to multiply again

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Daddyleo you forgot to multiply again \(\color{#0cbb34}{\text{End of Quote}}\) Multiply again by what number

Daddyleo:

read

Daddyleo:

den multiply aging the number you see previously

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Daddyleo den multiply aging the number you see previously \(\color{#0cbb34}{\text{End of Quote}}\) 6/3 =2 ×20 = 40 -5 = 35 of what ?

jhonyy9:

of 35 is 5

Daddyleo:

35x5

jhonyy9:

and what about this 175 ?

jhonyy9:

that's all ?

jhonyy9:

is this a math game ?

prettygirl09:

lol

Daddyleo:

Lol mah came up with it

Daddyleo:

Lol u smart people

prettygirl09:

me tired lol